# Donghyeon Lee1, Minbyul Jeong2∗, Jaewoo Kang1,3

### v1

@@ Abstract
Large language models (LLMs) have brought significant changes to many aspects of our lives. However, assessing and ensuring their chronological knowledge remains challenging. Existing approaches fall short in addressing the accumulative nature of knowledge, often relying on a single time stamp. To overcome this, we introduce ChroKnowBench, a benchmark dataset designed to evaluate chronologically accumulated knowledge across three key aspects: multiple domains, time dependency, temporal state. Our benchmark distinguishes between knowledge that evolves (e.g., scientific discoveries, amended laws) and knowledge that remain constant (e.g., mathematical truths, commonsense facts). Building on this benchmark, we present ChroKnowledge (Chronological Categorization of Knowledge), a novel sampling-based framework for evaluating and updating LLMs’ non-parametric chronological knowledge. Our evaluation led to the following observations: (1) The ability of eliciting temporal knowledge varies depending on the data format that model was trained on. (2) LLMs partially recall knowledge or show a cut-off at temporal boundaries rather than recalling all aspects of knowledge correctly. Thus, we apply our ChroKnowPrompt, an in-depth prompting to elicit chronological knowledge by traversing step-by-step through the surrounding time spans. We observe that our framework successfully updates the overall knowledge across the entire timeline in both the biomedical domain (+11.9%) and the general domain (+2.8%), highlighting its positive effect in refining temporal knowledge. This non-parametric approach also enables knowledge updates not only in open-source models but also in proprietary LLMs, ensuring comprehensive applicability across model types. We perform a comprehensive analysis based on temporal characteristics of ChroKnowPrompt and validate the potential of various models to elicit intrinsic temporal knowledge through our method. [^1] Our datasets and code are publicly available at https://github.com/dmis-lab/ChroKnowledge
@@ 1 Introduction
Figure 1: The overview of time variant dataset generation in ChroKnowBench. We accumulate knowledge in three key aspects: (1) multiple domains: general, biomedical, legal, commonsense, and mathematics; (2) time dependency: as time goes by, changeable knowledge; (3) temporal state: dynamic (has evolved over period) and static (no change occurred during period).
@@
When we examine the issue closely, it’s not just a matter of hallucination but also about whether the alignment of knowledge, particularly regarding dates, is accurate. Ensuring that LLMs maintain current and contextually relevant knowledge over time is crucial and researchers have explored various ways to investigate and verify the knowledge within these models (Zhang et al., 2024a). Pioneering works investigate whether language model has an ability of knowledge base or not in diverse domains (Petroni et al., 2019; Sung et al., 2021). Many subsequent studies analyze how LLMs define knowledge (Dai et al., 2022; Mishra et al., 2024); exploit how LLMs represent their knowledge (Geva et al., 2023; Zheng et al., 2024) with temporal context (Kasai et al., 2023; Fatemi et al., 2024); edit the misleading aspects of knowledge (Manakul et al., 2023; Wang et al., 2024c).
@@
Here, we raise a question: “Do these methods sufficiently address the accumulative nature of knowledge?”. Current temporal-related approaches for evaluating and updating LLMs often focus on single time stamps, struggling to address the accumulative characteristics (Jang et al., 2022a; Ge et al., 2024), which are especially important in specialized domains such as scientific discoveries and amended laws. This limitation can lead to outdated or incomplete information, undermining the models’ effectiveness and safety. Addressing these challenges requires a comprehensive approach is needed—one that accounts for both the temporal evolution and the accumulative nature of knowledge.
@@
Thus, we introduce ChroKnowBench, a benchmark dataset designed to evaluate chronologically accumulated knowledge across three dimensions: time dependency (time variant and time invariant), multiple domains (general, biomedical etc.), and temporal state (dynamic and static). ChroKnowBench differentiates between knowledge that is subject to evolution (e.g., transfer situation of a soccer player, scientific discoveries, and amended legal regulations) and knowledge that remains invariant (e.g., mathematical truths and commonsense facts). We then classify domains based on whether they are influenced by the flow of time, considering the domain specificity. Finally, we set the time frame to categorize the knowledge as either changeable or steady (Section 3).
@@
Building on this benchmark, we also present ChroKnowledge (Chronological Categorization of Knowledge), a novel framework for assessing and enhancing the non-parametric chronological knowledge of LLMs. So, we start from analyzing how current open-source and proprietary LLMs work. As we expected, the time invariant knowledge shows steady in all time frame. However, the domain-specific characteristics significantly influence the representation of temporal knowledge from LLMs. More stable domains exhibit consistent performance, while more variable domains show more fluctuations. These observations highlight that we need a comprehensive approach to enhance representing temporal knowledge from LLMs (Section 5).
@@
To this end, our ChroKnowPrompt approach utilizes an in-depth chronological prompting strategy that traverses and updates knowledge across adjacent time spans, effectively addressing issues of partial recall and temporal boundaries (Section 6). In knowledge recall, our evaluation reveals significant improvements in the biomedical domain (+11.9%) and modest gains in the general domain (+2.8%). Our non-parametric approach allows for direct updates to both proprietary and open-source LLMs without extensive retraining, highlighting its practicality and broad applicability (Section 7). Our work emphasizes the importance of temporal context in eliciting LLMs knowledge, introducing a novel framework for updating their chronological knowledge. We perform a comprehensive analysis based on temporal characteristics of ChroKnowPrompt and validate the potential of various models to elicit intrinsic temporal knowledge through our method.
@@ 2.2 Eliciting Knowledge using diverse Templates
Since models prefer different formats when eliciting their knowledge, it is important to use varied approaches to accurately assess their understanding (Zhou et al., ). While we initially evaluate the model’s knowledge using a standard triplet format, relying on a single template may not sufficiently capture the full extent of the model’s knowledge. Thus, following Hendrycks et al. (2021), we also employ a well-known format, multiple-choice question answering (MCQA), to elicit the model’s knowledge more effectively. As a result, we propose two templates for measuring how much knowledge the model holds: triplets (hereafter referred to as Generation) and MCQA. Each template is designed with appropriate few-shot exemplars and corresponding matching rules. For example, in Generation, due to the complexity of evaluating responses, we apply fuzzy matching techniques to compare the generated responses against predefined labels. See Appendix A.1 for further details of few-shot exemplars, fuzzy matching rules, and examples of two templates.
@@ 3.1 Task definition
Our primary focus is a time variant knowledge across three domains (general, biomedical, and legal) with comparisons to time invariant knowledge across two domains (commonsense and mathematics). What knowledge would be the difference between time variant and invariant? The time variant knowledge has a specific object changing across a stream of time. For example, “Cristiano Ronaldo (s) was a member of sports team of (r) Manchester United F.C. ($o_{1}$) in 2009 ($t_{1}$) and Real Madrid CF ($o_{k}$) in 2018 ($t_{k}$)”. Likewise, we identify subject and object alias for each relation, then gather yearly changed objects. After accumulating object lists $o_{1},o_{2}$, …, $o_{m}$}, we de-duplicate and fill out the missing data in specific years based on available data; objects between Manchester United F.C. ($o_{1}$) in 2009 ($t_{1}$) and Real Madrid CF ($o_{k}$) in 2018 ($t_{k}$), missing objects between 2010 ($t_{2}$) to 2017 ($t_{k-1}$) filled with existing object of 2009 ($t_{1}$). The statistic of ChroKnowBench dataset is in Table 2.
@@ 3.2 Dataset Generation
To construct dataset, we select annual knowledge sources for each domain, possible to be aligned with each elements even though the corpus does not specifically mention about time stamp. For sources with structured triplets, we identify temporal affect-able relations that typically change over time, such as “position held”. As time variant knowledge refers to the knowledge that has the potential to change over time, we divide it into two temporal states for more fine grained results: (1) dynamic, where the knowledge has evolved over the accumulated period. (2) static, where no change occurred during the accumulated period though it has potential to be changed. Following the methodology outlined in Section 3.1, we track changes in objects to build the dynamic dataset, employing normalization and de-duplication for verification. Each object is checked with strict exact string match, then add into objects pool. Simultaneously, we identified unchanged objects over the same time frame to construct the static dataset. At the end, all data elements consist with an associated object pool $\{o_{1},o_{2},\ldots,o_{m}\}$} over time frames $\{t_{1},t_{2},\ldots,t_{m}\}$}.
Table 2: Statistics of our benchmark dataset. We categorize whether knowledge changes over time (Time Variant) or remains constant (Time Invariant). We provide five domains to measure the knowledge from LMs. We set the temporal state with dynamic (knowledge that changes within the time frame we have set) and static (knowledge that do not change within the time frame we have set).
@@
Time Variant general 8 (s, r, o, t) dynamic 8,330 Wikidata
@@
biomedical 14 (s, r, o, t) dynamic 7,345 UMLS
@@ 3.3 Time Variant & Invariant Knowledge
We sourced time variant knowledge from the general, biomedical, and legal domains. In general domain, we utilize Wikidata (Vrandečić & Krötzsch, 2014) dump to track object changes among the time frame using suggested time quantifiers. Collecting similar amounts of dynamic and static instances across eight relations, the result is formatted as $\{{s,r,o,t\}}$ quadruplet for each object and accompanied time stamp. For biomedical domain, we parse Unified Medical Language System (UMLS) (Bodenreider, 2004) metathesaurus data, where suggest yearly updated research in the domain, following previous work of BIOLAMA (Sung et al., 2021). Due to the slow pace of change in biomedical research, object pools in this domain shows slight expansions or contractions over time frame. In the legal domain, we employ the Code of Federal Regulations (CFR) (U.S. Government, ) to track regulatory changes, as they suggest collection and accumulation of change in regulations at the end of year. Starting from pre-processing unstructured xml data, we adopt a QA-like format with placeholder for object, tracked among time frame which ends to dynamic or static whether it change or not. Time invariant knowledge, which remains constant regardless of temporal context, is drawn from common-sense and mathematical domains. We process the CSKG (Ilievski et al., 2021) dataset of commonsense knowledge, and Math-KG (Wang, 2022) for covering areas like data structures and pure mathematics. Further details of our source data are provided in Appendix A.2.
@@ 4 Experimental Setup
We conduct the experiments on a variety of proprietary and open-source LLMs. Each model utilizes either an instruction-tuned or chat version to enhance instruction following during sampling. We enumerate the eight open-source and one proprietary LLMs: Llama-3.1-8B-Instruct (Meta, 2024), Llama-3-8B-Instruct (Dubey et al., 2024), Llama-2-7b-chat-hf (Touvron et al., 2023b), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3.5-mini-instruct (Abdin et al., 2024), SOLAR-10.7B-Instruct-v1.0 (Kim et al., 2023), gemma-2-9b-it (Team et al., 2024b), gemma-7b-it (Team et al., 2024a), and GPT-4o mini (OpenAI, 2024a). We focus on analyzing trends in the chronological knowledge captured by suggested models from various companies. These models differ in corpus coverage reflecting their respective cut-off dates. Details of our inference setups are in Appendix A.3.
@@ 5 ChroKnowledge: Chronological Categorization of Knowledge
Figure 2: Performance analysis of general domain. The left two figures show the tendency of dynamic knowledge in temporal results, with more fluctuations in recent knowledge compared to static knowledge, which is more stable but still shows slight variation across the plots. The template-wise results reveal a trend of minimal reliance on internal representation (top figures). Meanwhile, the MCQA templates are influenced by the model’s specialized capabilities, even managing to overcome the training cutoff in recent years (bottom figures).
In this section, we introduce ChroKnowledge (Chronological Categorization of Knowledge), a sampling-based framework designed to evaluate and update LLMs’ non-parametric chronological knowledge. Our methodology assesses the temporal capabilities of LLMs using two distinct templates, detailed in Section 2.2, and explores how current LLMs encapsulate temporal information.
@@ 5.1 Results of Representing Knowledge from Large Language Models
For testing model’s knowledge within the categorization, we sample five times for each knowledge to elicit it as possible in dynamic and static dataset. We present our findings across three different aspects: temporal-wise, template-wise, and domain-wise results. In Figure 2 and 3, we depict the results for all domains and especially address general and biomedical domains in main section. Legal and invariant datasets are presented in Appendix (Figure 6 and 7).
@@
By comparing the left and right side would provide the tendency of temporal-wise results in Figure 2 and 3. A common trend across models is a decline in recent knowledge, aligning with pretraining corpus cutoff dates. In dynamic datasets, models show reduced knowledge in multiple-choice question answering (MCQA) settings, highlighting the impact of formatting and prompting on temporal abilities. Static datasets exhibit less fluctuation, though the result of general domain suggests domain-specific characteristics that each relations are more influenced via temporal sensitivity.
@@
By comparing the top and bottom side would provide the tendency of template-wise results in Figure 2 and 3. As we utilized two templates: generation and MCQA, the result is also checked in template specificity. Generation templates reveal a greater decline in recent knowledge, as models rely on internal representations without predefined answers. In contrast, MCQA templates help models select correct answers from options, mitigating some gaps in recent knowledge. Notably, GPT-4o Mini performs best in generation tasks, while gemma-2-9b-it excels in MCQA. In addition, the result of Phi-3.5-mini-instruct indicates specialized abilities can offset smaller model sizes, as it is more than twice small billion size compared to other models. This highlights the importance of template design in eliciting temporal knowledge, with generation benefiting from larger models and MCQA leveraging structured formats to enhance knowledge recall.
Figure 3: Performance analysis of biomedical domain. Across all the cases, there is little variability, highlighting a domain-specific feature. Both the left and right figures show similar aspects, though the failure to track updates between 2022 and 2023 is more clearly shown in the dynamic dataset. When comparing the top two generation figures with the MCQA results at the bottom, the generation templates show a narrower gap between models, while the gap in MCQA is more distinct. However, in terms of temporal states, the generation settings reveal clearer distinctions between dynamic and static knowledge, whereas in MCQA, the task-specific abilities may compensate for the differences.
@@
Comparing the Figure 2 and 3 provide the tendency of domain-wise results. In general domain, the models show a decline in recent knowledge, especially with generation templates. It comes from domain’s nature, such as changes in relation ‘position held’ or ‘member of sports team’ which has sensitivity to temporal cues, leading to higher variability. However, in biomedical domain, models fail to track recent updates between 2022 and 2023, particularly in generation tasks. Static datasets remain consistent, reflecting the stable nature of scientific knowledge. Here, MCQA setting offers some resilience though still limited to specific model’s improvement. We provide the overall trends of legal domain and time-invariant knowledge in Appendix A.4.
Overall, domain-specific characteristics significantly influence LLMs’ temporal knowledge representation. More stable domains like biomedical and legal exhibit consistent performance, while more variable domains like general show more fluctuations. These insights underscore the need for tailored strategies to enhance LLMs’ temporal knowledge capabilities.

### v2

@@ Abstract
Large language models (LLMs) have brought significant changes to many aspects of our lives. However, assessing and ensuring their chronological knowledge remains challenging. Existing approaches fall short in addressing the temporal adaptability of knowledge, often relying on a fixed time-point view. To overcome this, we introduce ChroKnowBench, a benchmark dataset designed to evaluate chronologically accumulated knowledge across three key aspects: multiple domains, time dependency, temporal state. Our benchmark distinguishes between knowledge that evolves (e.g., personal history, scientific discoveries, amended laws) and knowledge that remain constant (e.g., mathematical truths, commonsense facts). Building on this benchmark, we present ChroKnowledge (Chronological Categorization of Knowledge), a novel sampling-based framework for evaluating LLMs’ non-parametric chronological knowledge. Our evaluation led to the following observations: (1) The ability of eliciting temporal knowledge varies depending on the data format that model was trained on. (2) LLMs partially recall knowledge or show a cut-off at temporal boundaries rather than recalling all aspects of knowledge correctly. Thus, we apply our ChroKnowPrompt, an in-depth prompting to elicit chronological knowledge by traversing step-by-step through the surrounding time spans. We observe that it successfully recalls objects across both open-source and proprietary LLMs, demonstrating versatility, though it faces challenges with dynamic datasets and unstructured formats. [^1] Our datasets and code are publicly available at https://github.com/dmis-lab/ChroKnowledge
@@ 1 Introduction
Figure 1: The overview of ChroKnowBench. We gather knowledge with time stamps and separate them in three key aspects: (1) multiple domains: general, biomedical, legal, commonsense, and mathematics; (2) time dependency: as time goes by, changeable knowledge or not; (3) temporal state: dynamic (has evolved over period) and static (no change occurred during period).
@@
When we examine the issue closely, it’s not just a matter of hallucination but also about whether the alignment of knowledge, particularly regarding dates, is accurate. Ensuring that LLMs maintain current and contextually relevant knowledge over time is crucial and researchers have explored various ways to investigate and verify the knowledge within these models (Zhang et al., 2024b). Pioneering works investigate whether language model has an ability of knowledge base or not in diverse domains (Petroni et al., 2019; Sung et al., 2021). Many subsequent studies analyze how LLMs define knowledge (Dai et al., 2022; Mishra et al., 2024); exploit how LLMs represent their knowledge (Geva et al., 2023; Zheng et al., 2024) with temporal context (Kasai et al., 2023; Fatemi et al., 2024); edit the misleading aspects of knowledge (Manakul et al., 2023; Wang et al., 2024c).
@@
Here, we raise a question: “Do these methods sufficiently address the temporal adaptability of knowledge?”. Current temporal-related approaches for evaluating and updating LLMs often focus on single time stamps, struggling to address the adaptable characteristics of knowledge over time (Jang et al., 2022a; Ge et al., 2024), which are especially important in specialized domains such as scientific discoveries and amended laws. This limitation can lead to outdated or incomplete information, undermining the models’ effectiveness and safety. Addressing these challenges requires a comprehensive approach—temporal evolution, a core component of the knowledge accumulation.
@@
Thus, we introduce ChroKnowBench, a benchmark dataset designed to evaluate chronologically accumulated knowledge across three dimensions: time dependency (time variant and time invariant), multiple domains (general, biomedical etc.), and temporal state (dynamic and static). ChroKnowBench differentiates between knowledge that is subject to evolution (e.g., transfer situation of a soccer player, scientific discoveries, and amended legal regulations)—focusing on transformations in object-specific attributes such as roles or affiliations while keeping the subject and relation fixed—and knowledge that remains invariant (e.g., mathematical truths and commonsense facts). This object-level focus allows for precise and interpretable assessments of temporal knowledge dynamics by isolating changes in object-specific attributes. We then classify domains based on whether they are influenced by the flow of time, considering the domain specificity. Finally, we set the time frame to categorize the knowledge as either changeable or steady (Section 3).
@@
Building on this benchmark, we also present ChroKnowledge (Chronological Categorization of Knowledge), a novel framework for assessing and enhancing the non-parametric chronological knowledge of LLMs. So, we start from analyzing how current open-source and proprietary LLMs work. As we expected, the time invariant knowledge shows steady in all time frame. However, for time variant dataset, the domain-specific characteristics significantly influence the representation of temporal knowledge from LLMs. More stable domains exhibit consistent performance, while more variable domains show more fluctuations. These observations highlight that we need a comprehensive approach to enhance representing temporal knowledge from LLMs (Section 5).
@@
To this end, our ChroKnowPrompt approach utilizes an in-depth chronological prompting strategy that traverses knowledge across adjacent time spans, effectively addressing issues of partial recall and temporal boundaries (Section 6). In knowledge recall, our evaluation reveals improvements in the biomedical domain (11.2%) and the general domain (2.6%), shifting knowledge category from Partial Known to Known for unchanged objects. Our non-parametric approach allows for direct updates to both proprietary and open-source LLMs without extensive retraining, highlighting practicality and broad applicability (Section 7). Our work emphasizes the importance of temporal context in eliciting LLMs knowledge while identifying challenges with dynamic datasets and unstructured, context-rich formats like in the legal domain. A comprehensive analysis advocates for integrating parametric techniques to complement prompting and achieve more robust temporal knowledge handling.
@@ 2.2 Eliciting Knowledge using diverse Templates
Since models prefer different formats when eliciting their knowledge, it is important to use varied approaches to accurately assess their understanding (Zhou et al., ). While we initially evaluate the model’s knowledge using a standard triplet format, relying on a single template may not sufficiently capture the full extent of the model’s knowledge. Thus, following Hendrycks et al. (2021); Huang et al. (2024), we also employ a well-known format, multiple-choice question answering (MCQA) with 4 options, and True/False to elicit the model’s knowledge more effectively. As a result, we propose three templates for measuring how much knowledge the model holds: triplets (hereafter referred to as Generation), MCQA, and TF. Each template is designed with appropriate few-shot exemplars and corresponding matching rules. For example, in Generation, due to the complexity of evaluating responses, we apply fuzzy matching techniques to compare the generated responses against predefined labels. See Appendix A.2 for further details of few-shot exemplars, fuzzy matching rules, and examples of three templates.
@@ 3.1 Task definition
Our primary focus is a time variant knowledge across three domains (general, biomedical, and legal) with comparisons to time invariant knowledge across two domains (commonsense and mathematics). What knowledge would be the difference between time variant and invariant? The time variant knowledge has a specific object changing across a stream of time. For example, “Cristiano Ronaldo (s) was a member of sports team of (r) Manchester United F.C. ($o_{1}$) in 2009 ($t_{1}$) and Real Madrid CF ($o_{k}$) in 2018 ($t_{k}$)”. Particularly, we adopts an object-level focus, emphasizing changes in object attributes such as roles or affiliations. This approach ensures scalability by simplifying temporal dynamics, precision by capturing nuanced updates, and flexibility by supporting fine-grained real-world transformations without rigid relational constraints. Likewise, we identify subject and object alias for each relation, then gather yearly changed objects. After accumulating object lists $o_{1},o_{2}$, …, $o_{m}$}, we de-duplicate and fill out the missing data in specific years based on available data; objects between Manchester United F.C. ($o_{1}$) in 2009 ($t_{1}$) and Real Madrid CF ($o_{k}$) in 2018 ($t_{k}$), missing objects between 2010 ($t_{2}$) to 2017 ($t_{k-1}$) filled with existing object of 2009 ($t_{1}$). The statistic of ChroKnowBench dataset is in Table 2, and detail of object-level focus is in Appendix A.3.2.
@@ 3.2 Dataset Generation
To construct dataset, we select annual knowledge sources for each domain, possible to be aligned with each elements even though the corpus does not specifically mention about time stamp. For sources with structured triplets, we identify temporal affect-able relations that typically change over time, such as “position held”. As time variant knowledge refers to the knowledge that has the potential to change over time, we divide it into two temporal states for more fine grained results: (1) dynamic, where the knowledge has evolved over the period. (2) static, where no change occurred during the period, though it has potential to be changed. Following the methodology outlined in Section 3.1, we track changes in objects to build the dynamic dataset, employing normalization and de-duplication for verification. Each object is checked with strict exact string match, then add into objects pool. Simultaneously, we identified unchanged objects over the same time frame to construct the static dataset. At the end, all data elements consist with an associated object pool $\{o_{1},o_{2},\ldots,o_{m}\}$} over time frames $\{t_{1},t_{2},\ldots,t_{m}\}$}.
Table 2: Statistics of our benchmark dataset. We categorize whether knowledge changes over time (Time Variant) or remains constant (Time Invariant). Among five domains, we set the temporal state with dynamic (knowledge that changes within the time frame we have set) and static (knowledge that do not change within the time frame we have set). The number in parentheses represents the average change in objects per element within a dynamic dataset. See details in Appendix A.3.1.
@@
Time Variant general 8 (s, r, o, t) dynamic (2.6) 8,330 Wikidata
@@
biomedical 14 (s, r, o, t) dynamic (2.3) 7,345 UMLS
@@ 3.3 Time Variant & Invariant Knowledge
We sourced time variant knowledge from the general, biomedical, and legal domains. In general domain, we utilize Wikidata (Vrandečić & Krötzsch, 2014) dump to track object changes among the time frame using suggested time quantifiers. Collecting similar amounts of dynamic and static instances across eight relations, the result is formatted as $\{{s,r,o,t\}}$ quadruplet for each object and accompanied time stamp. For biomedical domain, we parse Unified Medical Language System (UMLS) (Bodenreider, 2004) metathesaurus data, where suggest yearly updated research in the domain, following previous work of BIOLAMA (Sung et al., 2021). Due to the slow pace of change in biomedical research, object pools in this domain shows slight expansions or contractions over time frame. In the legal domain, we employ the Code of Federal Regulations (CFR) (U.S. Government, ) to track regulatory changes, as they suggest collection and accumulation of change in regulations at the end of year. Starting from pre-processing unstructured xml data, we adopt a QA-like format with placeholder for object, tracked among time frame which ends to dynamic or static whether it change or not. Time invariant knowledge, which remains constant regardless of temporal context, is drawn from common-sense and mathematical domains. We process the CSKG (Ilievski et al., 2021) dataset of commonsense knowledge, and Math-KG (Wang, 2022) for covering areas like data structures and pure mathematics. Further details are provided in Appendix A.3, especially the sources and the mechanism of object-level focus. Especially, compared to TKGs (Zhang et al., 2024a), which focus on temporal snapshots, our ChroKnowBench emphasizes the detailed temporal evolution of individual knowledge elements, offering better adaptability for gradual changes; see Appendix A.3.3.
Figure 2: Performance analysis of general domain. (A) Heatmap in Generation template. For both dynamic and static datasets, a common trend across models is that performance is stronger in the intermediate years but decline recent years, reflecting the data-cutoff point. Dynamic knowledge shows more variation compared to static. Full results of total time frame is in Figure 10. (B) Template-wise performance for selected years. As time goes by, performance in generation goes low, on the other hand, MCQA and TF appeal to be rising. (C) Distribution of object changes in dynamic dataset.
@@ 4 Experimental Setup
We conduct the experiments on a variety of proprietary and open-source LLMs. Each model utilizes either an instruction-tuned or chat version to enhance instruction following during sampling. We enumerate the eight open-source and two proprietary LLMs: Llama-3.1-8B-Instruct (Meta, 2024), Llama-3-8B-Instruct (Dubey et al., 2024), Llama-2-7b-chat-hf (Touvron et al., 2023b), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3.5-mini-instruct (Abdin et al., 2024), SOLAR-10.7B-Instruct-v1.0 (Kim et al., 2023), gemma-2-9b-it (Team et al., 2024b), gemma-7b-it (Team et al., 2024a) for major open-source models, and GPT-4o mini (OpenAI, 2024a), Gemini-1.5-flash (DeepMind, 2024) for proprietary models. We focus on anlayzing trends in the chronological knowledge captured by those models, differ in corpus coverage. Details of our inference setups are in Appendix A.4.
@@ 5 ChroKnowledge: Chronological Categorization of Knowledge
In this section, we introduce ChroKnowledge (Chronological Categorization of Knowledge), a sampling-based framework designed to evaluate LLMs’ non-parametric chronological knowledge. Our methodology assesses the temporal capabilities of LLMs using two distinct templates, detailed in Section 2.2, and explores how current LLMs encapsulate temporal information.
Figure 3: Performance analysis of biomedical domain. The format of figure is same as Figure 2. (A) Compared to the general domain, both dynamic and static datasets show lower variability, reflecting a domain-specific tendency toward consistency in knowledge changes. Both of them shows performance decrease between 2022 and 2023, aligning with the cutoff pattern noted in the general domain. (B) As time goes by, performance in generation declines, but MCQA and TF continue to perform well.
@@ 5.1 Results of Representing Knowledge from Large Language Models
For testing model’s knowledge within the categorization, we sample five times for each knowledge to elicit it as possible in dynamic and static dataset. We present our findings across three different aspects: temporal-wise, template-wise, and domain-wise results. In Figure 2, 3 and 4, we depict the results for all time variant domains; general, biomedical and legal domains in main section. Time invariant datasets are presented in Appendix A.5, Figure 9.
@@
Comparing the upper and lower panels of (A) in Figure 2, 3 and 4 provides the tendency of temporal-wise results based on the generation results. A common trend across models is a decline in performance on recent knowledge, reflecting the pretraining corpus cutoff dates. Particularly in dynamic datasets, models demonstrate strong performance on earlier knowledge but experience a steeper decline in later periods, particularly in both general and biomedical domains. In contrast, static datasets show less fluctuation, with more stable yet weaker performance, highlighting limited temporal sensitivity due to reliance on a single timestamp. These results emphasize the need for frequent model updates, especially for dynamic knowledge, to ensure temporal relevance.
@@
(B) of Figure 2, 3 and 4 provide the average scores of template-wise results, for more template specificity checking in three templates: generation, MCQA, and TF. Generation templates reveal a greater decline in recent knowledge, as models rely on internal information without predefined answers. In contrast, MCQA and TF templates help models select correct answers from structured options and predefined formats, mitigating some gaps in recent knowledge. This trend is more evident in the biomedical domain with MCQA templates and the legal domain with TF, revealing how domain-specific knowledge is more effectively elicited through specific formats. Also, dynamic dataset in the general domain is more sensitive to temporal shifts than static, highlighting the importance of task-specific templates in eliciting and improving temporal robustness.
@@
Comparing the Figure 2, 3 and 4 provide the tendency of domain-wise results, demonstrating distinct domain-specific characteristics of temporal knowledge change. In general domain, the models show a decline in recent knowledge, with a more unstable distribution of scores, which stems from domain’s nature; changes in relations ‘position held’ or ‘member of sports team’ are more sensitive to temporal cues, leading to higher variability. Here, MCQA setting offers some resilience as it mitigates the knowledge decline observed in generation templates across dynamic and static datasets over different years. Every domain shows similar distribution of object changes with benchmark statistics, comparing (C) in each figure with Figure 8. This indicates that the models perform robustly as intended by the benchmark, without bias toward specific low changed objects. We provide more details for legal domain and time-invariant knowledge in Appendix A.5.
Overall, domain-specific characteristics significantly influence LLMs’ temporal knowledge representation. More stable domains like biomedical and legal exhibit consistent performance with time invariant knowledge, while general domain shows more inconsistency. These insights underscore the need for tailored strategies to enhance LLMs’ temporal knowledge capabilities.
Figure 4: Performance analysis of legal domain. The format of figure is same as Figure 2. (A) Among time variant domains, legal domain shows the most stable results of static, while the gap between dynamic and static datasets is the largest among domains. (B) When it comes to each template, generation shows the lowest performance, while TF settings perform extraordinarily well in answering correctly. (C) In the legal domain, the distribution shows the lowest number of object changes over time, supporting the conclusion of the stable results in the heatmap.